The length of a side of an equilateral triangle is represented by 2x-1. If the perimeter of the triangle is 21, what is the value of x ?
Answer by mananth(16946)

All sides are equal
perimeter = 3(2x-1)
3(2x-1)=21
2x-1= 21/3
2x-1=7
2x=7+1
2x=8
x=8/2
x=4